OpenStudy (aliaaalihassan):

Assume the function has an inverse. Without solving for the inverse find the indicated function values: f(x)=x^3+4x-1, (a f^-1 (4)

OpenStudy (perl):

you can use definition of inverse if (x,y) is in f , then (y,x) is in f^-1

OpenStudy (aliaaalihassan):

I've a problem with factorizing x^3+4x-5=0

OpenStudy (gorv):

no need to factor

OpenStudy (perl):

we have by definition f^-1 (y)=x <=> f(x) = y Therefore in our example f^-1 (4) = x <=> f(x) = 4 now solve f(x) = 4

OpenStudy (perl):

f(x) = 4 x^3+4x-1 = 4 x^3 + 4x - 5 = 0 (easy solution x = 1 ) so f(1) = 4 so f^-1 (4) = 1

OpenStudy (aliaaalihassan):

My problem is from where did the x=1 came from! i need a way to factorize x^3 + 4x - 5 = 0 to give me x=1 other than trying numbers that satisfy the equation mentally